# Deconstructing the Goldilocks Zone of Neural Network Initialization

### v1

@@ 6 Conclusion
This paper studies the Goldilocks zone of neural network initializations in the context of homogeneous architectures, both analytically and empirically. We demonstrate that the positive curvature of the training loss requires a robust top eigenspace of the Hessian inherited from its positive semi-definite component, the G-term and not a particular initialization norm per se. Normally, the G-term dominates in the Gauss-Newton decomposition of the Hessian but vanishes, for example, at extreme initialization scales or softmax temperatures, which is formally shown in Equation 5. We study properties of the G-term itself and relate high positive curvature to low model confidence, low initial cross-entropy loss, and a previously unknown type of vanishing gradient caused by the match between class priors found in the data and computed by the model. Finally, we report the training behavior of homogeneous networks optimized by gradient descent from a wide variety of initialization norms and uncover interesting training modalities at the edges of the Goldilocks zone (Figure 7). Furthermore, we find that model performance is only vaguely related to the Goldilocks zone. For example, homogeneous networks sometimes develop zero logits for an increasing number of inputs even when initialized and trained within the Goldilocks zone; thus, we argue against using positive curvature of the loss Hessian as a reliable estimate of the model performance.
@@
Most of our analysis requires neural networks to be homogeneous. Individual examinations are warranted for architectures with residual connections or non-homogeneous activation functions. Interestingly, our preliminary experiments show that BatchNorm eliminates the Goldilocks zone altogether, showing no excess of positive curvature even for normally initialized networks. Furthermore, while we comprehensively studied the behaviors of scaled homogeneous models when optimized by gradient descent, the conclusions might change with adaptive algorithms such as AdaGrad or Adam as these optimizers are known to effectively escape flat loss regions (Orvieto et al., 2022). We note that existing studies propose logit normalization for calibration not only after but also during training to improve convergence and generalization (Wei et al., 2022; Agarwala et al., 2023). Given the connection between model confidence and the Goldilocks zone unveiled in our study, future research may investigate how these methods affect positive curvature or, more generally, the eigenspectrum of the loss Hessian. Finally, our findings demonstrate the need for alternative measures to evaluate the quality of network initializations as positive curvature fails to faithfully capture model performance after training.
@@ Acknowledgements
The authors thank Mahalakshmi Sabanayagam, Jonathan Niles-Weed, and Kyunghyun Cho for numerous enlightening discussions. AV and JK were supported by the NSF Award 1922658. The Flatiron Institute is a division of the Simons Foundation. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.
@@ Appendix A Hessian Decomposition: Related Works
The Gauss-Newton decomposition in Equation 4 is a common entry point for many studies on the Hessian of large neural networks. The Hessian exhibits a “bulk-outlier” eigenspectrum with the majority of eigenvalues small and clustered around zero and only a handful of large positive outliers (Sagun et al., 2016; Gur-Ari et al., 2018; Ghorbani et al., 2019). This decomposition is inherited from the individual spectra of $G_{*}$ and $H_{*}$ with the top and the bulk eigenvalues attributed to these two terms, respectively. A few observations offer an intuitive explanation for this claim.
@@
First, the G-term can be rewritten in matrix form as $G_{*}(p)=J^{\top}[\text{diag}(p)-pp^{\top}]J$ where $J$ is the logit Jacobian. The matrix $\text{diag}(p)-pp^{\top}$ is positive semi-definite, so $G_{*}$ has non-negative eigenvalues. At the same time, $G_{*}$ computed on a single training sample is at most rank $K-1$, suggesting that the full G-term should be low-rank, especially for overparameterized architectures and small training datasets (Pennington & Bahri, 2017). Moreover, Fort & Ganguli (2019) find empirically that logit gradients on same-class training samples considerably overlap, which implies that eigenspaces of the G-terms of different data samples align. In support of this claim, Papyan (2019) shows that the top eigenvectors of the Hessian are attributable to high magnitude gradient class means. This agrees with a common observation that the Hessian has $K$ outlier eigenvalues. Recent studies argue that the corresponding eigenvectors encode the decision boundary of the network, and that more outliers emerge for more complex decision boundaries (e.g., after training from an adversarial initialization) (Liu et al., 2020; Sabanayagam et al., 2023). This observation can potentially correspond with a three-level hierarchical decomposition of the G-term demonstrated by Papyan (2020), which reveals additional clusters of smaller outliers.
@@
The H-term, on the other hand, has a continuous bulk of small eigenvalues, as independently verified by a number of studies (Papyan, 2018). Assuming that $H_{*}$ is a Wigner random matrix, Pennington & Bahri (2017) find its limiting spectral density to be consistent with experiments on randomly-initialized ReLU networks without biases. They also note that the H-term has a block off-diagonal structure for ReLU networks, which in particular implies that its positive curvature is zero. In the context of Neural Tangent Kernel (NTK), a kernel-based model for optimization dynamics, logits depend linearly on model parameters, and so $H_{*}$ is identically zero (Jacot et al., 2018).
@@
This pervasive analytical and empirical evidence suggests that, normally, the G-term has a pronounced bulk+outlier eigenspectrum that dominates the bulk eigenvalues found in the H-term. Pennington & Bahri (2017) note that the eigenspaces of these two matrices are in a nearly generic position and do not align in any special way. To this end, spiked matrix theory suggests that the Hessian eigenspectrum should be well approximated by the combination of individual eigenspectra of $G_{*}$ and $H_{*}$ (Perry et al., 2018). Since outlier eigenvalues of the Hessian are particularly important, researchers often drop the H-term from Equation 4 and focus on the G-term alone (Sagun et al., 2016; Fort & Ganguli, 2019; Cohen et al., 2021; Wang et al., 2022).
Appendix B Eigenvalues of $\mathbb{E}G_{*}$
@@
$\lambda_{i}(\mathbb{E}_{e}G_{*})=\begin{cases}(1-\lVert p\rVert^{2})\sigma_{e}^{2}+d\sigma_{c}^{2}\tilde{\lambda}_{i}&\text{ if }i\leq K-1,\\ (1-\lVert p\rVert^{2})\sigma_{e}^{2}&\text{ otherwise. }\end{cases}$
where $\tilde{\lambda}_{i}$ is the $i$-th largest eigenvalue of the matrix $\text{diag}(p)-pp^{\top}$. Since this is a rank-$1$ update to a diagonal matrix, its eigenvalues are interlaced with those of $\text{diag}(p)$; more precisely, for all $i\in[K-1]$, $p_{k_{i}}\geq\tilde{\lambda}_{i}\geq p_{k_{i+1}}$ where index $\{k_{i}\}_{i=1}^{K}$ sorts elements of $p$ in non-increasing order (Bunch et al., 1978). This, for example, immediately proves that when $p$ is uniform, all $K-1$ outlying eigenvalues of $\mathbb{E}_{e}G_{*}$ are identically $1/K$ as they get squeezed between values $p_{k}$. The same “bi-level” eigenstructure emerges when the expected G-term is taken across multiple training samples, as long as their softmax output remains uniform, as is the case for completely unconfident models ($\alpha\ll 1$).
@@ Appendix C The Limiting Behavior of $\mathbf{\text{diag}(p)-pp^{\top}}$
Positive curvature $\mathbf{\Gamma_{p}}$.
In Section 4, we defined $\Gamma_{p}$ to be positive curvature of the matrix $\text{diag}(p)-pp^{\top}$. By inspecting the elements of this matrix, we find
@@
Figure 10: The pre-collapse regime of LeNet-300-100 on FashionMNIST. Top row: the top eigenvector of the first layer’s Hessian is reshaped into the weight matrix dimensions ($300\times 784$), averaged across the hidden dimension ($300$), and reshaped as an input image ($28\times 28$). Note that the above heatmaps have different scales. As we gently increase $\alpha$, only one input sample (a picture of a bag in this case) remains uncertain, taking full control of the G-term and representing the principle curvature direction for the input layer. Bottom: The collapse of per-sample softmax entropies $\mathbb{H}(p^{\mu})$ and the corresponding positive curvature of the loss Hessian.
@@ Appendix D The Pre-collapse Regime
Having examined positive curvature of low-confident models in Section 4, we focus our analysis on the other end of the Goldilocks zone associated with high model confidence. Recall that positive curvature of a positive semi-definite matrix is lower bounded by $1$. Figure 2 illustrates this property for the G-term: as we increase $\alpha$, $\text{Tr}(G_{*})/\lVert G_{*}\rVert_{F}$ approaches $1$, remains level for a short time (we call this the pre-collapse regime), and then suddenly drops to 00. This property even transfers to the full Hessian, as seen from Figure 2. Further, in the pre-collapse regime of fully-connected architectures, we mysteriously observe that the principle curvature direction of the first layer’s Hessian aligns with one particular sample from the dataset (Figure 10). It turns out that we have all the necessary tools to explain this surprising behavior. As we increase $\alpha$, the G-terms $G_{*}^{\mu}$ of individual samples approach zero because $\text{diag}(p^{\mu})-p^{\mu}{p^{\mu}}^{\top}$ does; however, they do so at different rates. When $\alpha$ falls within a very specific narrow range, there is only one sample $\mu_{0}$ (the least confidently classified one) left with a large enough G-term, so that $G_{*}=\frac{1}{N}\sum_{\mu}^{N}G_{*}^{\mu}\approx G_{*}^{\mu_{0}}$. At this point, the top Hessian eigenspace comes from that of $G_{*}^{\mu_{0}}$ alone, which is spanned by the corresponding logit gradients. Since logit gradients with respect to parameters in the input layer are just the input vector itself, it manifests as the top eigenvector of the Hessian. Furthermore, in Appendix C, we show that $G_{*}(p)$ has only one non-zero eigenvalue when $p$ is on a knife edge from collapsing to a one-hot vector, so that positive curvature of $G^{\mu_{0}}$ and, hence, of the full Hessian, approaches $1$ in the pre-collapse regime.

### v2

@@ 6 Conclusion
This paper studies the Goldilocks zone of neural network initializations in the context of homogeneous architectures, both analytically and empirically. We demonstrate that the excess of positive curvature of the training loss requires a robust top eigenspace of the Hessian inherited from its positive semi-definite component, the G-term and not a particular initialization norm per se. Normally, the G-term dominates in the Gauss-Newton decomposition of the Hessian but vanishes, for example, at extreme initialization scales or softmax temperatures, which is formally shown in Equation 5. We study properties of the G-term itself and relate high positive curvature to low model confidence, low initial cross-entropy loss, and a previously unknown type of vanishing gradient caused by the match between class priors found in the data and computed by the model. Finally, we report the training behavior of homogeneous networks optimized by gradient descent from a wide variety of initialization norms and uncover interesting training modalities at the edges of the Goldilocks zone (Figure 7). Furthermore, we find that the connection between successful model training and the Goldilocks zone needs reconsideration. For example, homogeneous networks sometimes develop zero logits for an increasing number of inputs even when initialized and trained within the Goldilocks zone; thus, we argue against using positive curvature of the loss as a reliable estimate of the model performance.
@@
The scope of our study is limited to homogeneous neural networks. At the same time, the methodology developed in our study is applicable to inhomogeneous models and does not require scaling of the initialization that could potentially introduce unexpected and difficult to analyze degeneracies to the model (e.g., saturation of the activation functions such as TanH or Sigmoid). Recall that our analyses in Section 3 indicates that the Goldilocks zone can be exactly replicated by adjusting the logit temperature $T$ instead of $\alpha$-scaling the initialization, see Figure 1 and Equation 5. Thus, in our study of the Goldilocks zone and positive curvature, $\alpha$-scaling of the initialization norm can be equivalently replaced by varying $T$. This does not imply that our conclusions for homogeneous architectures transfer to inhomogeneous ones. Instead, our findings in Section 3 allow for repeating the analyses and experiments regarding the Goldilocks zone presented in Sections 4 and 5 for any type of model without having to scale the initialization.
Even though we scope our study at the initialization stage to conform with previous research, much of our analysis is more generally applicable at any point of the parameter space, including any points along training. All arguments in Section 3 remain valid because the Hessian exhibits the "bulk+outlier" eigenspectrum throughout optimization (in fact, as Papyan (2020) shows, it becomes even more pronounced with training) inherited from the H- and G-terms, respectively. All properties derived in Section 4 that do not explicitly assume random weights (e.g., initial loss) hold more generally beyond initialization because the adopted random model developed by Fort & Ganguli (2019) does. In their paper, they empirically confirmed that logit gradients cluster around their corresponding means not only at initialization, but also throughout training and at the solution.
In addition, while we comprehensively studied the behaviors of scaled homogeneous models when optimized by gradient descent, the conclusions might change with adaptive algorithms such as AdaGrad or Adam as these optimizers are known to effectively escape flat loss regions (Orvieto et al., 2022). We note that existing studies propose logit normalization for calibration not only after but also during training to improve convergence and generalization (Wei et al., 2022; Agarwala et al., 2023). Given the connection between model confidence and the Goldilocks zone unveiled in our study, future research may investigate how these methods affect positive curvature or, more generally, the eigenspectrum of the loss Hessian. Finally, our findings demonstrate the need for alternative measures to evaluate the quality of network initializations as positive curvature fails to faithfully capture model performance after training.
@@ Acknowledgements
The authors thank Mahalakshmi Sabanayagam, Jonathan Niles-Weed, and Kyunghyun Cho for numerous enlightening discussions. AV and JK acknowledge support through NSF Award 1922658. The Flatiron Institute is a division of the Simons Foundation. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Hessian Decomposition: Related Works
The Gauss-Newton decomposition in Equation 4 is a common entry point for many studies on the Hessian of large neural networks. The Hessian exhibits a “bulk-outlier” eigenspectrum with the majority of eigenvalues small and clustered around zero and only a handful of large positive outliers (Sagun et al., 2016; Gur-Ari et al., 2018; Ghorbani et al., 2019). This decomposition is inherited from the individual spectra of $\mathcal{G}$ and $\mathcal{H}$ with the top and the bulk eigenvalues attributed to these two terms, respectively. A few observations offer an intuitive explanation for this claim.
@@
First, the G-term can be rewritten in matrix form as $\mathcal{G}(p)=J^{\top}[\text{diag}(p)-pp^{\top}]J$ where $J$ is the logit Jacobian. The matrix $\text{diag}(p)-pp^{\top}$ is positive semi-definite, so $\mathcal{G}$ has non-negative eigenvalues. At the same time, $\mathcal{G}$ computed on a single training sample is at most rank $K-1$, suggesting that the full G-term should be low-rank, especially for overparameterized architectures and small training datasets (Pennington & Bahri, 2017). Moreover, Fort & Ganguli (2019) find empirically that logit gradients on same-class training samples considerably overlap, which implies that eigenspaces of the G-terms of different data samples align. In support of this claim, Papyan (2019b) shows that the top eigenvectors of the Hessian are attributable to high magnitude gradient class means. This agrees with a common observation that the Hessian has $K$ outlier eigenvalues. Recent studies argue that the corresponding eigenvectors encode the decision boundary of the network, and that more outliers emerge for more complex decision boundaries (e.g., after training from an adversarial initialization) (Liu et al., 2020; Sabanayagam et al., 2023). This observation can potentially correspond with a three-level hierarchical decomposition of the G-term demonstrated by Papyan (2020), which reveals additional clusters of smaller outliers.
@@
The H-term, on the other hand, has a continuous bulk of small eigenvalues, as independently verified by a number of studies (Papyan, 2019a). Assuming that $\mathcal{H}$ is a Wigner random matrix, Pennington & Bahri (2017) find its limiting spectral density to be consistent with experiments on randomly-initialized ReLU networks without biases. They also note that the H-term has a block off-diagonal structure for ReLU networks, which in particular implies that it has zero excess of positive curvature. In the context of Neural Tangent Kernel (NTK), a kernel-based model for optimization dynamics, logits depend linearly on model parameters, and so $\mathcal{H}$ is identically zero (Jacot et al., 2018).
@@
This pervasive analytical and empirical evidence suggests that, normally, the G-term has a pronounced bulk+outlier eigenspectrum that dominates the bulk eigenvalues found in the H-term. Pennington & Bahri (2017) note that the eigenspaces of these two matrices are in a nearly generic position and do not align in any special way. To this end, spiked matrix theory suggests that the Hessian eigenspectrum should be well approximated by the combination of individual eigenspectra of $\mathcal{G}$ and $\mathcal{H}$ (Perry et al., 2018). Since outlier eigenvalues of the Hessian are particularly important, researchers often drop the H-term from Equation 4 and focus on the G-term alone (Sagun et al., 2016; Fort & Ganguli, 2019; Cohen et al., 2021; Wang et al., 2022).
Appendix B Eigenvalues of $\mathbb{E}\mathcal{G}$
@@
$\lambda_{i}(\mathbb{E}_{e}\mathcal{G})=\begin{cases}(1-\lVert p\rVert^{2})\sigma_{e}^{2}+d\sigma_{c}^{2}\tilde{\lambda}_{i}&\text{ if }i\leq K-1,\\ (1-\lVert p\rVert^{2})\sigma_{e}^{2}&\text{ otherwise. }\end{cases}$
where $\tilde{\lambda}_{i}$ is the $i$-th largest eigenvalue of the matrix $\text{diag}(p)-pp^{\top}$. Since this is a rank-$1$ update to a diagonal matrix, its eigenvalues are interlaced with those of $\text{diag}(p)$; more precisely, for all $i\in[K-1]$, $p_{k_{i}}\geq\tilde{\lambda}_{i}\geq p_{k_{i+1}}$ where index $\{k_{i}\}_{i=1}^{K}$ sorts elements of $p$ in non-increasing order (Bunch et al., 1978). This, for example, immediately proves that when $p$ is uniform, all $K-1$ outlying eigenvalues of $\mathbb{E}_{e}\mathcal{G}$ are identically $1/K$ as they get squeezed between values $p_{k}$. The same “bi-level” eigenstructure emerges when the expected G-term is taken across multiple training samples, as long as their softmax output remains uniform, as is the case for completely unconfident models ($\alpha\ll 1$).
@@ Appendix C The Limiting Behavior of $\mathbf{\text{diag}(p)-pp^{\top}}$
Excess of positive eigenvalues $\mathbf{\Gamma_{p}}$.
In Section 4, we defined $\Gamma_{p}$ to be the excess of positive eigenvalues of the matrix $\text{diag}(p)-pp^{\top}$. By inspecting the elements of this matrix, we find
@@
Figure 10: The pre-collapse regime of LeNet-300-100 on FashionMNIST. Top row: the top eigenvector of the first layer’s Hessian is reshaped into the weight matrix dimensions ($300\times 784$), averaged across the hidden dimension ($300$), and reshaped as an input image ($28\times 28$). Note that the above heatmaps have different scales. As we gently increase $\alpha$, only one input sample (a picture of a bag in this case) remains uncertain, taking full control of the G-term and representing the principle curvature direction for the input layer. Bottom: The collapse of per-sample softmax entropies $\mathbb{H}(p^{\mu})$ and the corresponding positive curvature of the loss.
@@ Appendix D The Pre-collapse Regime
Having examined positive curvature of low-confident models in Section 4, we focus our analysis on the other end of the Goldilocks zone associated with high model confidence. Recall that excess of positive eigenvalues of a positive semi-definite matrix is lower bounded by $1$. Figure 2 illustrates this property for the G-term: as we increase $\alpha$, $\text{Tr}(\mathcal{G})/\lVert\mathcal{G}\rVert_{F}$ approaches $1$, remains level for a short time (we call this the pre-collapse regime), and then suddenly drops to 00. This property even transfers to the full Hessian, as seen from Figure 2. Further, in the pre-collapse regime of fully-connected architectures, we mysteriously observe that the principle curvature direction of the first layer’s Hessian aligns with one particular sample from the dataset (Figure 10). It turns out that we have all the necessary tools to explain this surprising behavior. As we increase $\alpha$, the G-terms $\mathcal{G}^{\mu}$ of individual samples approach zero because $\text{diag}(p^{\mu})-p^{\mu}{p^{\mu}}^{\top}$ does; however, they do so at different rates. When $\alpha$ falls within a very specific narrow range, there is only one sample $\mu_{0}$ (the least confidently classified one) left with a large enough G-term, so that $\mathcal{G}=\frac{1}{N}\sum_{\mu}^{N}\mathcal{G}^{\mu}\approx\mathcal{G}^{\mu_{0}}$. At this point, the top Hessian eigenspace comes from that of $\mathcal{G}^{\mu_{0}}$ alone, which is spanned by the corresponding logit gradients. Since logit gradients with respect to parameters in the input layer are just the input vector itself, it manifests as the top eigenvector of the Hessian. Furthermore, in Appendix C, we show that $\mathcal{G}(p)$ has only one non-zero eigenvalue when $p$ is on a knife edge from collapsing to a one-hot vector, so that excess of positive eiegnvalues of $\mathcal{G}^{\mu_{0}}$ and, hence, of the full Hessian, approaches $1$ in the pre-collapse regime.
@@ Appendix E More Connections to Optimization
Appendix F Alternative Estimation of Positive Curvature
Figure 13: Comparing methods for estimating positive curvature of the loss. Left: our method adopted from (Fort & Scherlis, 2019) Right: Hutchinson’s stochastic trace estimation.
As stated in Section 2, following the original work of Fort & Scherlis (2019), we estimate positive curvature of the loss by projecting the Hessian down onto a low-dimensional hyperplane using a sparse orthogonal transformation, and computing the quantity in Equation 2 explicitly for this smaller matrix. Alternatively, we could use the Hutchinson’s stochastic trace estimation method to approximate the trace and the norm of the original Hessian $H$ (Hutchinson, 1990). In this approach, we use Hessian-vector products to compute $\text{Tr}(H)\approx\frac{1}{m}\sum_{i=1}^{m}x_{i}^{\top}Hx_{i}$ and $\lVert H\rVert_{F}^{2}\approx\frac{1}{m}\sum_{i=1}^{m}(Hx_{i})^{\top}(Hx_{i})$ for $x_{i}\sim\{+1,-1\}^{P}$. As Figure 13 illustrates, Hutchinson’s method agrees with the one adopted in this study fairly well, capturing the boundaries and shape of the Goldilocks zone.